The stores sell ten different styles of yoga pants with identical unit costs and selling prices. A unit is defined as one pair of yoga pants. Each store has a manager who is paid a fixed salary. Individual salespeople receive a fixed salary and a sales commission. The store is trying to determine the desirability of opening another store, which is expected to have the following revenue and cost relationships:

Per Pair

Unit variable data:

Selling price

$60

Cost of pants

$29.50

Sales commissions

$4.50

Total variable costs

$34

Annual fixed costs:

Rent

$260,000

Salaries

$200,000

Advertising

$90,000

Other fixed costs

$70,000

Total fixed costs

$620,000

2. What is the annual breakeven point in (a) units sold and (b) revenues for the prospective store?

  1. If 55,000 units are sold, what will be the store’s operating profit (loss)?
  2. If sales commissions were discounted for individual salespeople in favor of an $41,000 increase in fixed salaries, what would be the annual breakeven point in (a) units sold and (b) revenues?
  3. Refer to the original data. If the store manager were paid $1.30 per unit sold in addition to her current fixed salary, what would be the annual breakeven point in (a) units sold and (b) revenues?
  4. Refer to the original data. If the store manager were paid $1.30 per unit commission on each unit sold in excess of the breakeven point, what would be the store’s operating profit if 70,000 units were sold? (This is $1.30 is in addition to both the commission paid to the sales staff and the store manager’s fixed salary.)
  5. Calculate the number of units sold where the operating profit under (a)a fixed salary plan and (b) a lower fixed salary and commission plan (for salespeople only) would be equal. Above that number of units sold, one plan would be more profitable than the other; below that number of units sold, the reverse would occur.
  6. Calculate the operating profit or loss under each plan in question 7 (above) at a sales level of (a) 50,000 units and (b) 60,000 units.
  7. Suppose the target operating profit is $368,000. How many units must be sold to reach the target under (a) the fixed salary plan and (b) the lower fixed salary and commission plan?
Answer #1

Dear Student,

As per the HOMEWORKLIB POLICY only the first four questions should be ansred. Kindly take note of it

Part 2

Break-even point (in units) = fixed costs / contribution margin per unit

Contribution margin per unit = selling price per unit – variable cost per unit

Contribution margin per unit = 60-34 =$26

Break-even point (in units) = 620000/26 = 23846 units

Break-even revenues = 23846 units * $60 = $1430760

Part A

Sales (60*55000)

3300000

Less: Variable costs (34*55000)

1870000

Contribution margin

1430000

Less: Fixed costs

620000

Operating income

$810000

Part B

Break-even point (in units) = fixed costs / contribution margin per unit

Contribution margin per unit = selling price per unit – variable cost per unit

Contribution margin per unit = 60-34 =$26

New fixed cost = 620000+41000 = 661000

Break-even point (in units) = 661000/26 = 25423 units

Break-even revenues = 25423 units * $60 = $1525380

Part C

Break-even point (in units) = fixed costs / contribution margin per unit

Contribution margin per unit = selling price per unit – variable cost per unit

New variable cost = 34+1.30 = 35.30

Contribution margin per unit = 60-35.30 =$24.70

Break-even point (in units) = 620000/24.70 = 25101 units

Break-even revenues = 25101 units * $60 = $1506060